# UIKA: Fast Universal Head Avatar from Pose-Free Images

### v1

@@ Abstract
We present UIKA, a feed-forward animatable Gaussian head model from an arbitrary number of unposed inputs, including a single image, multi-view captures, and smartphone-captured videos. Unlike the traditional avatar method, which requires a studio-level multi-view capture system and reconstructs a human-specific model through a long-time optimization process, we rethink the task through the lenses of model representation, network design, and data preparation. First, we introduce a UV-guided avatar modeling strategy, in which each input image is associated with a pixel-wise facial correspondence estimation. Such correspondence estimation allows us to reproject each valid pixel color from screen space to UV space, which is independent of camera pose and character expression. Furthermore, we design learnable UV tokens on which the attention mechanism can be applied at both the screen and UV levels. The learned UV tokens can be decoded into canonical Gaussian attributes using aggregated UV information from all input views. To train our large avatar model, we additionally prepare a large-scale, identity-rich synthetic training dataset. Our method significantly outperforms existing approaches in both monocular and multi-view settings. Project page: https://zijian-wu.github.io/uika-page/
@@ 1 Introduction
Table 1: Comparison with state-of-the-art 3D head reconstruction methods. $F$ denotes a feed-forward pipeline that requires no test-time optimization or fine-tuning. PF indicates pose-free input, i.e., camera and expression parameters are not required. RTA denotes real-time animatability ($\geq$ 30 FPS).
@@ 2.1 Generative 2D Head Avatar
Recent years have witnessed significant advances in image-based talking head synthesis, with most state-of-the-art approaches operating within a 2D generative framework [62, 63, 3, 15, 64, 21, 68, 75, 86, 43, 41]. Two families of generative models have been widely adopted in recent years. Early work leverages Generative Adversarial Networks (GANs) to learn motion-driven warping of a static reference image. To enable faithful transfer of facial expressions and dynamic motion, diverse motion descriptors are used, including 2D facial keypoints [85, 63], depth maps [26], and latent embeddings [3]. Diffusion models, renowned for their generative capability, have been adopted in several recent approaches, especially for one-shot talking-face generation. Benefiting from strong priors learned from large-scale image datasets, diffusion-based methods [74, 44, 71, 14] can drive stylized or out-of-domain faces. However, their inherent 2D modeling assumptions and the lack of 3D awareness cause them to struggle with large pose variations and consistent head-movement synthesis.
2.2 Optimize-based 3D Head Avatar
Optimization-based methods for 3D head avatars commonly employ explicit or implicit 3D representations, including meshes [23], NeRFs [18, 97, 76, 27, 49, 84], SDFs [93], point-based representations [94], and 3D Gaussians [72, 6, 79, 92], to reconstruct per-subject avatars. Among these, methods that rely on monocular video input [18, 97, 76, 72, 6, 88] can reconstruct faithful 3D head avatars for the observed viewpoints but struggle to generalize to novel views. Other approaches [27, 49, 84, 79, 92] exploit large-scale multi-view datasets [33, 47, 52, 80, 1] to learn expressive, generalizable priors over both geometry and appearance, achieving state-of-the-art visual realism. Despite their impressive reconstruction quality, the need for per-subject optimization limits their practicality and broader applicability. More recently, diffusion-based frameworks such as CAP4D [66] synthesize multi-view images from a single portrait to guide avatar reconstruction. While this strategy improves identity preservation across views, it still requires time-consuming optimization, hindering deployment in real-time or one-shot settings.
2.3 Feed Forward 3D Head Avatar
Feed-forward avatar reconstruction methods create realistic, animatable 3D head avatars from as few as one or a handful of input images by learning expressive priors from large-scale datasets, whether monocular [73] or multi-view [33, 47]. Prior work [49, 5, 83, 38, 39, 46, 95, 67, 8, 81, 82] introduces NeRFs [48] for feed-forward reconstruction of 3D head avatars, delivering high-fidelity results and precise control over novel camera viewpoints. To enhance performance, several methods incorporate 3D supervision either from monocular data [9, 11, 17] or synthetic multi-view data [87, 12, 13, 42]. However, NeRF-based architectures struggle to support real-time avatar animation. Recent methods [7, 25, 34] have demonstrated the utility of 3D Gaussian Splatting (3DGS) [31] for avatar modeling, achieving rapid rendering without sacrificing visual quality. Nevertheless, GAGAvatar [7] relies on a 2D-to-3D lifting scheme that fails to reconstruct regions not visible in the input plausibly. LAM [25] suffers reduced reconstruction accuracy under novel viewpoints due to its one-shot setting. Avat3R [34] departs from the one-shot setting, requiring multiple input views and rerunning the entire network inference to generate the 3D head Gaussian attributes for each new expression. In contrast, the proposed UIKA reconstructs realistic, animatable 3D head avatars from multiple unposed input images in a feed-forward manner.

### v2

@@ Abstract
We present UIKA, a feed-forward animatable Gaussian head model from an arbitrary number of unposed inputs, including a single image, multi-view captures, and smartphone-captured videos. Unlike the traditional avatar method, which requires a studio-level multi-view capture system and reconstructs a human-specific model through a long-time optimization process, we rethink the task through the lenses of model representation, network design, and data preparation. First, we introduce a UV-guided avatar modeling strategy, in which each input image is associated with a pixel-wise facial correspondence estimation. Such correspondence estimation allows us to reproject each valid pixel color from screen space to UV space, which is independent of camera pose and character expression. Furthermore, we design learnable UV tokens on which the attention mechanism can be applied at both the screen and UV levels. The learned UV tokens can be decoded into canonical Gaussian attributes using aggregated UV information from all input views. To train our large avatar model, we additionally prepare a large-scale, identity-rich synthetic training dataset. Our method significantly outperforms existing approaches in both monocular and multi-view settings. See more details in our project page: https://zijian-wu.github.io/uika-page/.
@@ 1 Introduction
Table 1: Comparison with state-of-the-art 3D head reconstruction methods. $F$ denotes a feed-forward pipeline that requires no test-time optimization or fine-tuning. PF indicates pose-free input, i.e., camera and expression labels are not required. RTA denotes real-time animatability ($\geq$ 30 FPS).
@@ 2.1 Generative 2D Head Avatar
Early one-shot 2D methods are typically built upon conditional GANs [17, 77, 21, 28, 29, 101, 71] and explicit warping [96, 67, 25, 3, 66, 68, 75, 11, 12], where a motion representation is estimated to deform a static reference image. More recently, diffusion models [83, 45, 79, 13] have substantially improved fidelity and robustness in one-shot portrait animation, benefiting from scalable architecture (e.g., DiT [56, 14, 51]) and strong priors learned from large-scale data. Diffusion-based talking-head synthesis typically falls into two lines: audio-driven models [70, 84, 94] that generate speech-conditioned facial motion with realistic articulation and landmark-controlled models [79, 45, 46] that use sparse keypoint constraints for controllable reenactment and cross-domain transfer. Despite rapid progress, 2D diffusion-based methods such as DiffusionRig [13] remain image-space with weakly 3D-aware performance, which often leads to limited extrapolation to extreme viewpoints, constrained long-range motion consistency, and high inference cost.
2.2 Optimization-based 3D Head Avatar
Optimization-based methods typically require long-time optimization for a specific person, spanning mesh-centric pipelines [22], volumetric neural fields [16, 26, 106, 50, 85, 95], and more recently animatable 3D Gaussians [87, 81, 6, 89, 103, 99]. Recently, generative models have been introduced as an auxiliary source of supervision. CAP4D [73] and MVP4D [72] leverage morphable multi-view diffusion to synthesize pseudo-observation supervision from reference images to regularize identity preservation and view completeness during reconstruction. Such hybrid pipelines narrow the gap between sparse-view inputs and multi-view reconstruction quality, but they rely on iterative fitting process and thus are still ill-suited for real-time or truly one-shot applications, motivating feed-forward 3D head avatar reconstruction.
2.3 Feed-Forward 3D Head Avatar
Feed-forward 3D head avatar methods orient to a single network forward pass reconstruction by learning strong identity priors from large-scale datasets, including in-the-wild monocular videos [82, 100], studio-quality multi-view captures [32, 48, 54, 90, 1] and synthetic data [10, 15, 98, 11, 12, 43]. Early approaches [5, 93, 38, 39, 47, 104, 74, 91, 92] can infer a personalized, animatable avatar from one or a handful of images. However, the NeRF-based [49] method, such as GPAvatar [8], struggles to support real-time avatar animation due to implicit modeling. Recent works [40, 80, 53, 57, 33] adopt 3D Gaussian Splatting [30] as a representation to enable real-time and high-fidelity rendering. For example, LAM [24] and GAGAvatar [7] can reconstruct a 3D head avatar through a single portrait input, but often degrade under large viewpoint extrapolation for unobserved regions. More recently, Avat3r [34] and HeadGAP [103] extend to multiple inputs to improve 3D consistency. However, they are trained totally on multi-view datasets with limited identity diversity, which can hinder generalization to in-the-wild inputs. FastAvatar [80] accepts an arbitrary number of inputs but suffers from redundant Gaussian points when using more and more views. In contrast, our method reconstructs high-fidelity and animatable head avatars in a single feed-forward pass from any number of images without requiring additional camera or expression annotations and supports joint training on identity-rich monocular videos and 3D-consistent multi-view data for efficient and scalable deployment.